# Arithmetic of cuts in ordered abelian groups and of ideals over valuation rings

### v4

@@ 1. Introduction
[…]
@@
In the present paper, we will exclusively work with the second definition, which we call the upper cut set addition. Every cut is uniquely determined by its upper cut set, so instead of cuts we will just work with final segments (for a reason that will be explained later). The set of final segments of $G$ is linearly ordered by inclusion, and unions, intersections and sums of any collections of final segments are again final segments.
@@
is the largest solution $T_{\rm max}$ for the inequality $S_{1}+T\subseteq S_{2}\,$; hence if a solution of (1) exists, then $S_{2}-_{\rm fs}S_{1}$ is the largest. However, in the literature there does not seem to exist any recipy for computing $S_{2}-_{\rm fs}S_{1}$ in terms of (elementwise) sums and differences of final segments derived from $S_{1}$ and $S_{2}$ that would help solve the problems that came up in the papers [3, 4] and which we will explain below.
@@
Take a valued field $(K,v)$ with value group $vK$ and valuation ring $\mathcal{O}_{v}$. For $a\in K$ we write $va$ in place of $v(a)$. When we talk of $\mathcal{O}_{v}$-ideals we will always assume them to be nonzero, and we also include fractional ideals; in other words, we talk about nonzero $\mathcal{O}_{v}$-modules $I\subsetneq K$. Just like the set of final segments of $vK$, the set of $\mathcal{O}_{v}$-ideals is linearly ordered by inclusion, and unions and intersections of any collections of $\mathcal{O}_{v}$-ideals are again $\mathcal{O}_{v}$-ideals. The function
@@
is the largest solution $J_{\rm max}$ for the inequality $I_{1}J\subseteq I_{2}\,$. Again, in the literature there does not seem to exist any recipy for computing $I_{1}:I_{2}$ from $I_{1}$ and $I_{2}$ that would help solve the problems addressed in Section 3.5. This section is devoted to the computation of the annihilators $I_{2}:I_{1}$ of quotients $I_{1}/I_{2}$ of $\mathcal{O}_{v}$-ideals $I_{1},I_{2}$ with $I_{2}\subseteq I_{1}\,$. The results are then applied in Section 3.6 to compute the annihilators of some special quotients that appear in [3] and [4] and to answer the question in which cases they are equal to the maximal ideal $\mathcal{M}_{v}$ of $\mathcal{O}_{v}\,$. Our answers to the above questions are given in Theorems 3.13 and 3.14.
@@ 2.1. Preliminaries on cuts and final segments
Take any ordered set $(\Gamma,\leq)$ (by “ordered”, we will always mean “totally ordered”). If $M_{1},M_{2}$ are nonempty subsets of $\Gamma$ and $\alpha\in\Gamma$, we will write $\alpha<M_{2}$ if $\alpha<\beta$ for all $\beta\in M_{2}$, and we will write $M_{1}<M_{2}$ if $\alpha<M_{2}$ for all $\alpha\in M_{1}$. Similarly, we use the relations $>$, $\leq$ and $\geq$ in place of $<$.
@@
For any subset nonempty $M\subseteq\Gamma$, we set
@@
That is, if $M^{-}$ is the smallest final segment of $\Gamma$ which contains $M$.
@@
3) If $\,\Gamma$ is densely ordered, then an element $\gamma\in G$ is an infimum of $S$ if and only if it is a supremum of $S^{c}$. ∎
Every ordered set $\Gamma$ carries a topology with the open intervals $(\alpha,\beta)$, $\alpha,\beta\in\Gamma$ with $\alpha<\beta$, as its basic open sets. We leave the proof of the following lemma to the reader.
@@
$\widehat{S}\>:=\>\left\{\begin{array}[]{ll}S\cup\{\gamma\}&\mbox{if $S$ has infimum $\gamma$ in $G$, and}\\ S&\mbox{otherwise.}\end{array}\right.$
@@ 2.2. Final segments in ordered abelian groups
If $S_{1}$ and $S_{2}$ are final segments, then so is $S_{1}+S_{2}\,$. With this addition, $G\uparrow$ is a commutative monoid with neutral element $0^{-}$. It is in general not a group and can even contain itempotents other than $0^{-}$; see part 1) of Lemma 2.3 and part 3) of Lemma 2.6. For example, the lexicographically ordered group $G=\mathbb{Z}\times\mathbb{Z}$ has the proper nontrivial subgroup $H=\{0\}\times\mathbb{Z}$, which gives rise to two idempotent final segments. Indeed, since $H+H=H$, both
$\{\alpha\in G\mid\forall\beta\in H:\alpha>\beta\})\quad\mbox{ and }\quad\{\alpha\in G\mid\exists\beta\in H:\alpha\geq\beta\})$
are idempotent, and it can be shown that together with $G$ itself they are the only idempotent final segments in $G$.
@@
[…]
@@
[…]
@@
4): Since $S$ is nonprincipal, part 3) yields that $S=S+0^{+}$. If $T$ is closed in $G$, then $\widehat{T}=T$ and there is nothing to show. If $T$ is closed in $G$, then by part 4) of Lemma 2.2, $T$ has an infimum $\alpha\in G\setminus T$. Thus, $\widehat{T}=\alpha^{-}$ and $0^{+}+\widehat{T}=\alpha^{+}=T$, so that $S+\widehat{T}=S+0^{+}+\widehat{T}=S+T$. ∎
@@ 2.3. Subtraction of final segments
The notion of subtraction of Dedekind cuts has been defined in articles in the literature, such as [6] and [9]. When upper cut set addition is used, then for two Dedekind cuts $\Lambda_{1}=(\Lambda_{1}^{L},\Lambda_{1}^{R})$ and $\Lambda_{2}=(\Lambda_{2}^{L},\Lambda_{2}^{R})$, their difference $\Lambda_{2}-\Lambda_{1}$ has been defined to be the cut with upper cut set $\Lambda_{2}^{R}-\Lambda_{1}^{L}$, which is the sum of the final segments $\Lambda_{2}^{R}$ and $-\Lambda_{1}^{L}$ and thus again a final segment.
@@
Note that $(-S)^{c}=-(S^{c})$ since $\gamma\in(-S)^{c}\Leftrightarrow\gamma>-S\Leftrightarrow-\gamma<S\Leftrightarrow-\gamma\in S^{c}\Leftrightarrow\gamma\in-(S^{c})$. Therefore we will omit the brackets and just write “$\,-S^{c}\,$”.
@@
[…]
In order to obtain a difference operation that extends the one of $G$, replacing “$>$” by $\geq$” in (18) we define the (modified) final segment subtraction of two final segments $S_{2}$ and $S_{1}$ as
@@
$\Delta\alpha^{+}\>=\>\left\{\begin{array}[]{ll}(-\alpha)^{-}&\mbox{if $G$ is densely ordered,}\\ (-\alpha-\gamma)^{-}&\mbox{if $G$ is discretely ordered.}\end{array}\right.$
@@
2) The sets $\Delta S_{1}$, $S_{2}-_{\rm mc}S_{1}\,$, $-S_{1}^{c}$ and $S_{2}-S_{1}^{c}$ are again final segments of $G$.
@@
2): For $\Delta S_{1}$ this holds by definition, so for $S_{2}-_{\rm mc}S_{1}$ it follows from part 1). For $-S_{1}^{c}$ we observe that $S_{1}^{c}$ is an initial segment, hence $-S_{1}^{c}$ is a final segment and the same holds for $S_{2}-S_{1}^{c}=S_{2}+(-S_{1}^{c})$.
@@
The question arises whether the difference operation $-_{\rm mc}$ provides some positive answer to our questions (QFS1) and (QFS2). As part 2) of Proposition 2.5 in conjunction with part 4) of Lemma 2.10 shows, it indeed gives a complete answer to these questions in the case where $S_{1}$ is principal. It now remains to deal with the case where $S_{1}$ is nonprincipal. This case is much more complex and we need some preparation.
@@ 2.4. Invariance groups
Take any non-empty subset $M\subseteq G$.
@@
1):  Take $\alpha,\beta\in{\mathcal{G}}(M)$. Then $M+\alpha+\beta=M+\beta=M$, whence $\alpha+\beta\in{\mathcal{G}}(M)$. Further, $M-\alpha=M+\alpha-\alpha=M$, whence $-\alpha\in{\mathcal{G}}(M)$. This proves that ${\mathcal{G}}(M)$ is a group.
@@
5):  By part 1), it suffices to show that ${\mathcal{G}}(M)$ is convex. Take $\alpha,\beta\in{\mathcal{G}}(M)$ and $\gamma\in G$ with $\alpha\leq\gamma\leq\beta$. Then for all $\delta\in M$, $\delta+\alpha\leq\delta+\gamma\leq\delta+\beta$ and $\delta+\alpha,\delta+\beta\in M$. Hence $\delta+\gamma\in M$ by convexity of $M$. This proves that $M+\gamma\subseteq M$. In the same way, we also obtain that $M-\gamma\subseteq M$ because $-\alpha,-\beta\in{\mathcal{G}}(M)$ (since ${\mathcal{G}}(M)$ is a group) and $-\beta\leq-\gamma\leq-\alpha$. By part 2) it follows that $\gamma\in{\mathcal{G}}(M)$.
@@
$\alpha+S\subseteq S\>\Leftrightarrow\>\alpha\in{\mathcal{G}}(S)^{-}\;\mbox{ and }\;\alpha+S\subsetneq S\Leftrightarrow\alpha\in{\mathcal{G}}(S)^{+}\>.$
@@
5):  We have ${\mathcal{G}}(\alpha+H)=\{\gamma\in G\mid\gamma+\alpha+H=\alpha+H\}=\{\gamma\in G\mid\gamma+H=H\}=H$. By equation (17), $(\alpha+H)^{-}=\alpha^{-}+H$. Hence by part 4) and equation (25) of Lemma 2.12, ${\mathcal{G}}((\alpha+H)^{-})={\mathcal{G}}(\alpha^{-}+H)={\mathcal{G}}(\alpha^{-})+H=H$. We observe that the complement of $-(\alpha+H)^{+}$ in $G$ is $(-\alpha+H)^{-}$. Applying part 3) of Lemma 2.12 twice, we obtain that ${\mathcal{G}}((\alpha+H)^{+})={\mathcal{G}}(-(\alpha+H)^{+})={\mathcal{G}}((-\alpha+H)^{-})=H$.
@@
Take final segments $S_{1}\,$ $S_{2}$ and $T$ such that $S_{1}+T=S_{2}$. Then ${\mathcal{G}}(S_{1})\subseteq{\mathcal{G}}(S_{1})\cup{\mathcal{G}}(T)={\mathcal{G}}(S_{2})$, and if ${\mathcal{G}}(S_{1})\subsetneq{\mathcal{G}}(S_{2})$, then ${\mathcal{G}}(T)={\mathcal{G}}(S_{2})$.
@@
1): The first and second equality are special cases of part 3) of Lemma 2.12. The equality ${\mathcal{G}}(S)={\mathcal{G}}(\Delta S)$ is seen as follows. If $S$ is nonprincipal, then by part 3) of Lemma 2.10, $\Delta S=-S^{c}$ and the equality follows from what we have already proved. If $S$ is principal, say $S=\alpha^{-}$ for some $\alpha\in G$, then $\Delta S=(-\alpha)^{-}$ by (20), whence ${\mathcal{G}}(\Delta S)=\{0\}={\mathcal{G}}(S)$ by (25).
@@
If $(\Lambda^{L},\Lambda^{R})$ is not a cut, then $\Lambda^{L}\cap\Lambda^{R}=\{\gamma\}$ for some $\gamma\in G$. It follows that $\Lambda^{R}=\gamma^{-}$, hence by part 1) of Lemma 2.16, ${\mathcal{G}}(\Lambda^{L})={\mathcal{G}}(\Lambda^{R})=\{0\}$. ∎
@@
If $\gamma=\alpha-\beta$ with $\alpha\in S$ and $\beta\in S^{c}$, then $\alpha-\gamma=\beta\notin S$, so $-\gamma\notin{\mathcal{G}}(S)$. Since ${\mathcal{G}}(S)$ is a subgroup of $G$, this implies that $\gamma\notin{\mathcal{G}}(S)$. Conversely, if $H$ is a subgroup of $G$ properly containing ${\mathcal{G}}(S)$, then there is $\gamma\in H^{>0}\setminus{\mathcal{G}}(S)$. Then $-\gamma\notin{\mathcal{G}}(S)$ and there is $\alpha\in S$ such that $\alpha-\gamma\notin S$. This implies $\beta=\alpha-\gamma\in S^{c}$, whence $\gamma=\alpha-\beta\in S-S^{c}$.
@@
1): This follows from equation (21) in part 3) of Lemma 2.10 together with Lemma 2.18. The equalities in (27) follow from part 1) together with parts 3) and 5) of Lemma 2.10. The inequality holds since ${\mathcal{G}}(S_{1})^{+}$ only contains positive elements.
@@
4) The set of all $\alpha\in G$ such that $\alpha+S\subseteq G^{>0}$ is equal to $-S^{c}$.
@@
1): If ${\mathcal{G}}(S)^{+}$ does not contain $S$, then $S\setminus{\mathcal{G}}(S)^{+}\neq\emptyset$, so there is some $\alpha\in S\cap{\mathcal{G}}(S)$. Since ${\mathcal{G}}(S)$ is a group, it follows that $-\alpha\in{\mathcal{G}}(S)$. Therefore $S=-\alpha+S$, whence $0=\alpha-\alpha\in S$. The converse holds since $0\in{\mathcal{G}}(S)^{+}$ is impossible.
@@
Now assume that ${\mathcal{G}}(S)^{+}=\gamma+S$ for some $\gamma\in G$. Then $-\gamma+{\mathcal{G}}(S)^{+}\subseteq S$, but $-\gamma\notin S$ since otherwise $0\in{\mathcal{G}}(S)^{+}$, which is impossible.

### v5

@@ Abstract.
We thank the referee for his many very helpful corrections and comments.
@@ 1. Introduction
Our approach is to translate questions about ideals to corresponding questions about final segments in the value group $vK$ of $(K,v)$. The latter are connected with the investigation of cuts in ordered abelian groups which has appeared at several points in the literature. For a survey and a list of references, see [8]. The present paper uses, and improves, parts of the unpublished manuscript [10]. Let us describe the background on cuts first.
@@
In the present paper, we will exclusively work with the second definition, which we call the upper cut set addition. Every cut is uniquely determined by its upper cut set, so instead of cuts we will just work with final segments (for a reason that will be explained later). The set of final segments of $G$ is linearly ordered by inclusion, and unions, intersections and sums of any collections of final segments are again final segments, as long as they are proper nonempty subsets of $G$.
@@
is the largest solution $T_{\rm max}$ for the inequality $S_{1}+T\subseteq S_{2}\,$; hence if a solution of (1) exists, then $S_{2}-_{\rm fs}S_{1}$ is the largest. However, in the literature there does not seem to exist any recipe for computing $S_{2}-_{\rm fs}S_{1}$ in terms of (elementwise) sums and differences of final segments derived from $S_{1}$ and $S_{2}$ that would help solve the problems that came up in the papers [3, 4] and which we will explain below.
@@
Take a valued field $(K,v)$ with value group $vK$ and valuation ring $\mathcal{O}_{v}$. For $a\in K$ we write $va$ in place of $v(a)$. When we talk of $\mathcal{O}_{v}$-ideals we will always assume them to be nonzero, and we also include fractional ideals; in other words, we talk about . Just like the set of final segments of $vK$, the set of $\mathcal{O}_{v}$-ideals is linearly ordered by inclusion, and unions, intersections and products of any collections of $\mathcal{O}_{v}$-ideals are again $\mathcal{O}_{v}$-ideals, as long as they are nonzero $\mathcal{O}_{v}$-modules properly contained in $K$. The function
@@
is the largest solution $J_{\rm max}$ for the inequality $I_{1}J\subseteq I_{2}\,$. Again, in the literature there does not seem to exist any recipe for computing $I_{1}:I_{2}$ from $I_{1}$ and $I_{2}$ that would help solve the problems addressed in Section 3.5. This section is devoted to the computation of the annihilators $I_{2}:I_{1}$ of quotients $I_{1}/I_{2}$ of $\mathcal{O}_{v}$-ideals $I_{1},I_{2}$ with $I_{2}\subseteq I_{1}\,$. The results are then applied in Section 3.6 to compute the annihilators of some special quotients that appear in [3] and [4] and to answer the question in which cases they are equal to the maximal ideal $\mathcal{M}_{v}$ of $\mathcal{O}_{v}\,$. Our answers to the above questions are given in Theorems 3.13 and 3.14.
@@ 2.1. Preliminaries on cuts and final segments
Take $(\Gamma,\leq)$ to be the underlying ordered set of any nonzero ordered abelian group (by “ordered”, we will always mean “totally ordered”). Note that $(\Gamma,\leq)$ does not have a smallest or largest element, and the definitions in this section also work without the presence of the group structure. If $M_{1},M_{2}$ are nonempty subsets of $\Gamma$ and $\alpha\in\Gamma$, we will write $\alpha<M_{2}$ if $\alpha<\beta$ for all $\beta\in M_{2}$, and we will write $M_{1}<M_{2}$ if $\alpha<M_{2}$ for all $\alpha\in M_{1}$. Similarly, we use the relations >>, $\leq$ and $\geq$ in place of <<.
@@
For any nonempty subset $M\subseteq\Gamma$, we set
@@
That is, if $M^{-}\neq\Gamma$, then it is the smallest final segment of $\Gamma$ which contains $M$.
@@
3) If $\Gamma$ is densely ordered, then an element $\gamma\in\Gamma$ is an infimum of $S$ if and only if it is a supremum of $S^{c}$. ∎
Every nonempty ordered set $(\Gamma,\leq)$ without smallest or largest element carries a topology with the open intervals $(\alpha,\beta)$, $\alpha,\beta\in\Gamma$ with $\alpha<\beta$, as its basic open sets. We leave the proof of the following lemma to the reader.
@@
$\widehat{S}\>:=\>\left\{\begin{array}[]{ll}S\cup\{\gamma\}&\mbox{if $S$ has infimum $\gamma$ in $\Gamma$, and}\\ S&\mbox{otherwise.}\end{array}\right$.
@@ 2.2. Final segments in ordered abelian groups
If $M$, $M_{1}$ and $M_{2}$ are nonempty subsets of $G$, then we set $M_{1}+M_{2}=\{\alpha+\beta\mid\alpha\in M_{1}\,,\,\beta\in M_{2}\}$ and $M_{1}-M_{2}=\{\alpha-\beta\mid\alpha\in M_{1}\,,\,\beta\in M_{2}\}$. From this, the sets $\alpha+M$, $\alpha-M$ and $M-\alpha$ are obtained via replacing $\alpha$ by $\{\alpha\}$. We set M:=0M-M:=0-M.
If $S_{1}$ and $S_{2}$ are final segments, then so is $S_{1}+S_{2}\,$. With this addition, $G\uparrow$ is a commutative monoid with neutral element $0^{-}$. It is in general not a group and can even contain itempotents other than $0^{-}$; see part 1) of Lemma 2.3 and part 3) of Lemma 2.6. For example, the lexicographically ordered group $G=\mathbb{Q}\times\mathbb{Z}$ has the proper nontrivial subgroup $H=\{0\}\times\mathbb{Z}$, which gives rise to two idempotent final segments. Indeed, since $H+H=H$, both
@@
are idempotent, and it can be shown that they are the only idempotent final segments in $G$.
@@
$S_{1}\,+\,S_{2}\>=\>\alpha_{1}^{-}\,+\,\alpha_{2}^{+}\>=\>\alpha_{1}^{-}\,+\,\alpha_{2}^{-}\,+\,0^{+}\>=\>(\alpha_{1}+\alpha_{2})^{-}\,+\,0^{+}\>=\>(\alpha_{1}+\alpha_{2})^{+}\>.$
@@
$S_{1}\,+\,S_{2}\>=\>\alpha_{1}^{+}\,+\,\alpha_{2}^{+}\>=\>\alpha_{1}^{-}\,+\,0^{+}\,+\,\alpha_{2}^{-}\,+\,0^{+}\>=\>(\alpha_{1}+\alpha_{2})^{-}\,+\,0^{+}\>=\>(\alpha_{1}+\alpha_{2})^{+}\>.$
@@
4): Since $S$ is nonprincipal, part 3) yields that $S=S+0^{+}$. If $T$ is closed in $G$, then $\widehat{T}=T$ and there is nothing to show. If $T$ is not closed in $G$, then by part 4) of Lemma 2.2, $T$ has an infimum $\alpha\in G\setminus T$. Thus, $\widehat{T}=\alpha^{-}$ and $0^{+}+\widehat{T}=\alpha^{+}=T$, so that $S+\widehat{T}=S+0^{+}+\widehat{T}=S+T$. ∎
@@ 2.3. Subtraction of final segments
The notion of subtraction of Dedekind cuts has been defined in articles in the literature, such as [6] and [9]. When upper cut set addition is used, then for two Dedekind cuts $\Lambda_{1}=(\Lambda_{1}^{L},\Lambda_{1}^{R})$ and $\Lambda_{2}=(\Lambda_{2}^{L},\Lambda_{2}^{R})$, their difference $\Lambda_{2}-\Lambda_{1}$ has been defined to be the cut with upper cut set $\Lambda_{2}^{R}-\Lambda_{1}^{L}$, which is the sum of the final segments $\Lambda_{2}^{R}$ and $\Lambda_{1}^{L}$ and thus again a final segment.
@@
Note that $(-S)^{c}=-(S^{c})$ since $\gamma\in(-S)^{c}\Leftrightarrow\gamma>-S\Leftrightarrow-\gamma<S\Leftrightarrow-\gamma\in S^{c}\Leftrightarrow\gamma\in-(S^{c})$. Therefore we will omit the brackets and just write $S^{c}\,$”.
@@
$\alpha^{-}\,-\,(\alpha^{-})^{c}\>=\>\alpha^{-}\,+\,\{\gamma\in G\mid\gamma>-\alpha^{-}\}\>=\>\alpha^{-}\,+\,\{\gamma\in G\mid\gamma>-\alpha\}\>=\>0^{+}\>.$
In order to obtain a difference operation that extends the one of $G$, replacing “>>” by $\geq$” in (18) we define the (modified) final segment subtraction of two final segments $S_{2}$ and $S_{1}$ as
@@
[…]
@@
2) The sets $\Delta S_{1}$, $S_{2}-_{\rm mc}S_{1}\,$, $S_{1}^{c}$ and $S_{2}-S_{1}^{c}$ are again final segments of $G$.
@@
2): For $\Delta S_{1}$ this holds by definition, so for $S_{2}-_{\rm mc}S_{1}$ it follows from part 1). For $S_{1}^{c}$ we observe that $S_{1}^{c}$ is an initial segment, hence $S_{1}^{c}$ is a final segment and the same holds for $S_{2}-S_{1}^{c}=S_{2}+(-S_{1}^{c})$.
@@
The question arises whether the difference operation $_{\rm mc}$ provides some positive answer to our questions (QFS1) and (QFS2). As part 2) of Proposition 2.5 in conjunction with part 4) of Lemma 2.10 shows, it indeed gives a complete answer to these questions in the case where $S_{1}$ is principal. It now remains to deal with the case where $S_{1}$ is nonprincipal. This case is much more complex and we need some preparation.
@@ 2.4. Invariance groups
Take any nonempty subset $M\subseteq G$.
@@
1): Take $\alpha,\beta\in{\mathcal{G}}(M)$. Then $M+\alpha+\beta=M+\beta=M$, whence $\alpha+\beta\in{\mathcal{G}}(M)$. Further, $M-\alpha=M+\alpha-\alpha=M$, whence $\alpha\in{\mathcal{G}}(M)$. This proves that ${\mathcal{G}}(M)$ is a group.
@@
5): By part 1), it suffices to show that ${\mathcal{G}}(M)$ is convex. Take $\alpha,\beta\in{\mathcal{G}}(M)$ and $\gamma\in G$ with $\alpha\leq\gamma\leq\beta$. Then for all $\delta\in M$, $\delta+\alpha\leq\delta+\gamma\leq\delta+\beta$ and $\delta+\alpha,\delta+\beta\in M$. Hence $\delta+\gamma\in M$ by convexity of $M$. This proves that $M+\gamma\subseteq M$. In the same way, we also obtain that $M-\gamma\subseteq M$ because $\alpha,-\beta\in{\mathcal{G}}(M)$ (since ${\mathcal{G}}(M)$ is a group) and $\beta\leq-\gamma\leq-\alpha$. By part 2) it follows that $\gamma\in{\mathcal{G}}(M)$.
@@
[…]
@@
5): We have ${\mathcal{G}}(\alpha+H)=\{\gamma\in G\mid\gamma+\alpha+H=\alpha+H\}=\{\gamma\in G\mid\gamma+H=H\}=H$. By equation (17), $(\alpha+H)^{-}=\alpha^{-}+H$. Hence by part 4) and equation (25) of Lemma 2.12, ${\mathcal{G}}((\alpha+H)^{-})={\mathcal{G}}(\alpha^{-}+H)={\mathcal{G}}(\alpha^{-})+H=H$. We observe that the complement of $(\alpha+H)^{+}$ in $G$ is $(-\alpha+H)^{-}$. Applying part 3) of Lemma 2.12 twice, we obtain that ${\mathcal{G}}((\alpha+H)^{+})={\mathcal{G}}(-(\alpha+H)^{+})={\mathcal{G}}((-\alpha+H)^{-})=H$.
@@
Take final segments $S_{1}\,$, $S_{2}$ and $T$ such that $S_{1}+T=S_{2}$. Then ${\mathcal{G}}(S_{1})\subseteq{\mathcal{G}}(S_{1})\cup{\mathcal{G}}(T)={\mathcal{G}}(S_{2})$, and if ${\mathcal{G}}(S_{1})\subsetneq{\mathcal{G}}(S_{2})$, then ${\mathcal{G}}(T)={\mathcal{G}}(S_{2})$.
@@
1): The first and second equality are special cases of part 3) of Lemma 2.12. The equality ${\mathcal{G}}(S)={\mathcal{G}}(\Delta S)$ is seen as follows. If $S$ is nonprincipal, then by part 2) of Lemma 2.8, $\Delta S=-S^{c}$ and the equality follows from what we have already proved. If $S$ is principal, say $S=\alpha^{-}$ for some $\alpha\in G$, then $\Delta S=(-\alpha)^{-}$ by (20), whence ${\mathcal{G}}(\Delta S)=\{0\}={\mathcal{G}}(S)$ by (25).
@@
If $(\Lambda^{L},\Lambda^{R})$ is not a cut, then $\Lambda^{L}\cap\Lambda^{R}=\{\gamma\}$ for some $\gamma\in G$. It follows that $\Lambda^{R}=\gamma^{-}$, hence by part 2) of Lemma 2.16, ${\mathcal{G}}(\Lambda^{L})={\mathcal{G}}(\Lambda^{R})=\{0\}$. ∎
@@
If $\gamma=\alpha-\beta$ with $\alpha\in S$ and $\beta\in S^{c}$, then $\alpha-\gamma=\beta\notin S$, so $\gamma\notin{\mathcal{G}}(S)$. Since ${\mathcal{G}}(S)$ is a subgroup of $G$, this implies that $\gamma\notin{\mathcal{G}}(S)$. Conversely, if $H$ is a subgroup of $G$ properly containing ${\mathcal{G}}(S)$, then there is $\gamma\in H^{>0}\setminus{\mathcal{G}}(S)$. Then $\gamma\notin{\mathcal{G}}(S)$ and there is $\alpha\in S$ such that $\alpha-\gamma\notin S$. This implies $\beta=\alpha-\gamma\in S^{c}$, whence $\gamma=\alpha-\beta\in S-S^{c}$.
@@
1): This follows from equation (21) in part 3) of Lemma 2.10 together with Lemma 2.18. The equalities in (27) follow from part 1) together with parts 3) and 5) of Lemma 2.10. The inclusion holds since ${\mathcal{G}}(S_{1})^{+}$ only contains positive elements.
@@
4) The set of all $\alpha\in G$ such that $\alpha+S\subseteq G^{>0}$ is equal to $S^{c}$.
@@
1): If ${\mathcal{G}}(S)^{+}$ does not contain $S$, then $S\setminus{\mathcal{G}}(S)^{+}\neq\emptyset$, so there is some $\alpha\in S\cap{\mathcal{G}}(S)$. Since ${\mathcal{G}}(S)$ is a group, it follows that $\alpha\in{\mathcal{G}}(S)$. Therefore $S=-\alpha+S$, whence $0=\alpha-\alpha\in S$. The converse holds since $0\in{\mathcal{G}}(S)^{+}$ is impossible.
@@
Now assume that ${\mathcal{G}}(S)^{+}=\gamma+S$ for some $\gamma\in G$. Then $\gamma+{\mathcal{G}}(S)^{+}\subseteq S$, but $\gamma\notin S$ since otherwise $0\in{\mathcal{G}}(S)^{+}$, which is impossible.